the they Also place in the class, to show that they work out with each problem. way should. your answers spaces provided (30 points) A propeller on an airplane is turned off when rotating at an angular speed of 4450 rev/min, and comes to rest after spinning through 12.510 rev 1. a. (15) What is the angular acceleration of the propeller blade, assumed to be constant? rad/s b. (15) How much time does it take before the propeller comes to rest?

This is question 1 please show all work calculus is used.
Answer #1

1. given wo = 4450 rev/min

comes to rest after theta = 12.510 rev

a. angular accelration of the propellor blade = alpha

then

wf = 0

wo = 4450 rev/min = 4450*2*pi rad/ 60 s= 466.0029 rad/s

theta = 12.510 rev = 12.510*2*pi rad = 78.6026 rad

hence

using equation

2*alpha*theta = wf^2 - wo^2

alpha = -1381.37099 rad/s/s

b. time taken for the propellor to come to rest = t

wf = wo + alpha*t

t = (wf - wo)/alpha = 0.337348115 s